# Conformal prediction after efficiency-oriented model selection

### v2

@@ Abstract
Given a family of pretrained models and a hold-out set, how can we construct a valid conformal prediction set while selecting a model that minimizes the width of the set? If we use the same hold-out data set both to select a model (the model that yields the smallest conformal prediction sets) and then to construct a conformal prediction set based on that selected model, we suffer a loss of coverage due to selection bias. Alternatively, we could further splitting the data to perform selection and calibration separately, but this comes at a steep cost if the size of the dataset is limited. In this paper, we address the challenge of constructing a valid prediction set after efficiency-oriented model selection. Our novel methods can be implemented efficiently and admit finite-sample validity guarantees without invoking additional sample-splitting. We show that our methods yield prediction sets with asymptotically optimal size under certain notion of continuity for the model class. The improved efficiency of the prediction sets constructed by our methods are further demonstrated through applications to synthetic datasets in various settings and a real data example.
@@ 1 Introduction
However, despite the success of the conformal prediction framework in offering statistical validity, the efficiency in the size of conformal prediction sets generally lacks a universal guarantee. Indeed, theories of the efficiency of conformal prediction have been established only for special cases—under specific assumptions on the data distribution and the corresponding regression algorithm used in implementing conformal prediction. For example, Sesia and Candès (2020) show that the conformalized quantile regression (Romano et al., 2019) achieves the optimal length asymptotically if the quantile regression function estimate is consistent and the data distribution has certain regularity properties. Similar results also hold for the split conformal method and the full conformal method when paired with the residual score (Lei et al., 2018). See also the work of Gyôrfi and Walk (2019); Burnaev and Vovk (2014); Sadinle et al. (2019); Izbicki et al. (2020), among others, for additional examples of settings where efficiency of conformal prediction has been established under specific conditions; in general, the efficiency of conformal prediction is closely tied to choosing a conformity score that is well suited to the data distribution, as has been studied empirically in Aleksandrova and Chertov (2021); Johansson et al. (2017); Sesia and Candès (2020).
Given the necessity of choosing the right model to construct an effective prediction set, and the practical reality that the most suitable model cannot be predetermined without analyzing the data, an integrated framework that combines model selection with the construction of prediction sets would be highly beneficial. This serves as the motivation of this paper. In this work, we address the challenge of constructing a finite-sample valid conformal prediction set while selecting a model that minimizes the size of the prediction set, from a given family of pretrained models. Specifically, fix some target coverage level $1-\alpha\in(0,1)$, e.g., 90%. Given a collection of pretrained models, a set of independent and identically distributed hold-out data $\mathcal{D}_{n}$, and a new test point $X_{n+1}$ from the same distribution, we would like to construct a conformal prediction set $\widehat{C}_{n}$ using a selected model (i.e., selected as the model that offers the smallest prediction interval), such that
@@
  • the size of the prediction set $\widehat{C}_{n}$ is as small as possible (i.e., we have benefited from selecting the most efficient model).
Several existing works in the literature have considered similar problems to the question described above. In Section 2.3 below, we will closely examine the work of Yang and Kuchibhotla (2024), which is most relevant to our work. In their work, Yang and Kuchibhotla (2024) propose either choosing the model with the shortest interval without correcting for selection to achieve efficiency (which suffers from selection bias unless the set of candidate models is extremely small), or using a sample splitting based approach (which removes the issue of selection bias, but suffers in statistical efficiency due to reduced sample size). In contrast, our work aims to correct for selection bias without any data splitting. We will compare our methods with theirs more closely below. For a more comprehensive discussion of related literature, we defer it to Section 6.1.
@@
We propose a novel method with two variants to construct finite-sample valid conformal prediction sets without additional sample-splitting after efficiency-oriented model selection. Both versions are computationally easy to implement, and offer validity with no assumptions aside from exchangeability of the data. In addition, we show that under certain continuity condition for the model class, our methods output prediction sets that are asymptotically optimal in size. We will see on simulated and real data that these methods offer improved finite-sample efficiency relative to existing approaches, by removing selection bias without paying the price of data splitting.
@@
We first formulate our problem as well as surveying existing methods in Section 2. We then present our proposed methods as well as coverage guarantees in Section 3. The asymptotic efficiency of our methods is investigated in Section 4. In Section 5, we conduct empirical comparisons across different methods to demonstrate the effectiveness of our methods in practice. Finally, we review the related literature and conclude with a brief discussion in Section 6.
@@ 2 Problem formulation
In this section, we first review some background on conformal prediction, which serves as the backbone of our prediction sets construction. We will then specify the efficiency-oriented selection procedure considered in this paper and formulate the problem of interest. Finally, we will briefly review the existing methods related to our problem.
@@ 2.1 Background: conformal prediction
Consider the learning problem of predicting the response $y\in\mathcal{Y}$ based on an observed feature $x\in\mathcal{X}$. The conformal prediction offers a framework of constructing valid prediction sets without placing any assumption on the data distribution $P$ other than data exchangeability.
@@
If the amount of available data is limited, we may wish to avoid the cost of data splitting inherent in split conformal prediction (that is, in split conformal prediction, $n$ calibration data points have been removed from the total quantity of available data, and thus the score function $S$ is trained using a smaller sample size.) The full conformal prediction (Vovk et al., 2005) provides a solution that allows us to run regression algorithm $\mathcal{A}$ and construct prediction set without data splitting.
Given a dataset $\mathcal{D}_{n}\in(\mathcal{X}\times\mathcal{Y})^{n}$ and an algorithm $\mathcal{A}:\cup_{n\geq 0}(\mathcal{X}\times\mathcal{Y})^{n}\to\{\textnormal{functions}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}\}$ that treats the input training data points symmetrically, the full conformal prediction set is defined as
$\widehat{C}_{\texttt{full-CP}}(X_{n+1})=\left\{y\in\mathcal{Y}:S^{y}(X_{n+1},y)\leq\mathrm{Quantile}_{(1-\alpha)(1+1/n)}\left(\left\{S^{y}(X_{i},Y_{i})\right\}_{i\in[n]}\right)\right\},$ (2)
@@
2.2 Model selection criterion: width of a prediction band
@@
We next need to specify how we choose one of these pretrained models—what does it mean to choose the smallest, or most efficient, model from the set? Formally, we will quantify model efficiency through a loss function $\mathcal{L}$, as follows:
@@
A typical choice for this loss might be the average size of the prediction set,
@@
where ||$\cdot$|| ⋅ | denotes the size of the set, e.g., the Lebesgue measure if $\mathcal{Y}=\mathbb{R}$ (i.e., the length, if $C_{q}^{\lambda}$ returns intervals), or the cardinality if $\mathcal{Y}$ is a finite set of categorical labels. However, this is not the only choice we might use in practice—for instance, we might choose median, rather than mean, size. [^2] As an example of another kind of loss function that we might use in a regression setting, the work of Liang et al. (2023) consider a selection procedure that aims at minimizing the squared estimation error, $\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-f_{\lambda}(X_{i}))^{2}$, rather than the interval length. Like our work which will be described in details below, Liang et al. (2023)’s approach uses the full conformal framework to then incorporate a hypothesize test point $(X_{n+1},y)$ into this selection process. Note however that this requires a loss $\mathcal{L}$ that depends on $(X,Y)$ values at evaluation (rather than only $X$ values), and is therefore not directly comparable to our framework.
@@
    the length of the interval (which is shared across all values $x$, and therefore does not depend on the particular evaluation points $x_{1},\dots,x_{m}$).
@@
Main aim: Given a collection of pretrained models $\{S^{\lambda}:\lambda\in\Lambda\}$, and a calibration data set $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$, we aim to output a prediction set at a new test point $X_{n+1}$ that is both valid (under exchangeability of the data) and efficient (in the sense of aiming to minimize $\mathcal{L}$ over the collection of models).
@@ 2.3 Existing methods: the work of Yang and Kuchibhotla (2024)
In their first approach, split conformal prediction sets are constructed for each pretrained model $S^{\lambda}$, and the one with the smallest size is selected directly without any correction for selection bias. We refer this approach to as YK-baseline in this paper, since it provides a baseline against which other methods (that do account for selection bias) can be compared. We can define it precisely as follows. First, for each $\lambda\in\Lambda$, define
@@
In other words, this is the split conformal prediction set constructed with the selected model $\hat{\lambda}\in\Lambda$—but without any correction for selection bias. [^3] Yang and Kuchibhotla (2024)’s work does not formally specify a mechanism for defining a “best” prediction set, so here we choose a specific definition for concreteness, defining “best” to mean achieving the minimal loss (e.g., minimal length) over the $n+1$ calibration and test feature values.
@@
By construction, YK-baseline yields a prediction set with the smallest size among all the possible split conformal prediction sets $C_{\hat{q}(\lambda)}^{\lambda}$ (indexed over $\lambda\in\Lambda$). If $\Lambda$ is large then we would expect to lose substantial coverage due to selection bias; at the other extreme, if $|\Lambda|=1$, then this is simply split conformal prediction (i.e., there is no selection since there is only a single possible choice of $\lambda$), and so marginal coverage will hold. More generally, for any number $|\Lambda|$ of candidate models, Yang and Kuchibhotla (2024) show the following concentration-based coverage guarantee:
@@
(here their result assumes the data points $(X_{1},Y_{1}),\dots,(X_{n},Y_{n}),(X_{n+1},Y_{n+1})$ to be i.i.d., rather than only assuming exchangeability). We can interpret this lower bound as saying that, if $|\Lambda|$ is not too large (i.e., $\log(2|\Lambda|)\ll n$), then the selection bias is limited and will not substantially reduce coverage. On the other hand, for large $|\Lambda|$, this lower bound does not give a meaningful guarantee (i.e., coverage may be arbitrarily low), as will be confirmed in our experiments in Section 5.
@@
To avoid the problem of selection bias incurred by YK-baseline, Yang and Kuchibhotla (2024) propose an alternative method, based on data splitting, which we will refer to as YK-split. At a high level, the calibration set of $n$ data points is split into two parts, where $n_{1}$ many data points are used to select a model, and the remaining $n_{2}=n-n_{1}$ data points are used to construct the split conformal prediction set with the selected model. Specifically, defining
@@ 3.1 The ModSel-CP method
As we will see in the proof, this marginal coverage result will follow by reformulating ModSel-CP as an instance of full conformal prediction. In particular, just as full conformal prediction may be run with any base algorithm $\mathcal{A}$ that maps data sets to fitted models, here the model selection step (i.e., the process of selecting a model $\hat{\lambda}(y)$ as a function of the data) plays the role of this algorithm $\mathcal{A}$.
@@ 3.1.1 Implementation
The ModSel-CP prediction set defined in (9) satisfies
@@
Furthermore, if it holds that:
@@
Note that Theorem 3.2 does not give an exact calculation for the prediction set $\widehat{C}_{\textnormal{{ModSel-CP}}}(X_{n+1})$, but rather provides upper and lower bounds on this set. In particular, we can equivalently rewrite the upper bound (11) as
@@
Hence, ModSel-CP (or rather, this upper bound on ModSel-CP) can be efficiently implemented as long as sets of the form $\left\{y:\mathcal{L}_{n+1}(\lambda,S^{\lambda}(X_{n+1},y))\leq q\right\}$ are easy to compute. For example, in the case of the residual score $S^{\lambda}(x,y)=|y-f_{\lambda}(x)|$, where $f_{\lambda}$ is a pretrained regression function, this set is equivalent to a union of intervals, i.e., the upper bound (11) is equivalent to
$\widehat{C}_{\texttt{ModSel-CP}}(X_{n+1})\subseteq\left\{y\in\mathbb{R}:\min_{\lambda\in\mathcal{M}}|y-f_{\lambda}(X_{n+1})|\leq\hat{q}(\hat{\lambda})\right\}=\bigcup\limits_{\lambda\in\mathcal{M}}\left[f_{\lambda}(X_{n+1})-\hat{q}(\hat{\lambda})\ ,\ f_{\lambda}(X_{n+1})+\hat{q}(\hat{\lambda})\right].$ (14)
In practice, then, we can think of the set defined in (11) as a more computationally efficient, but potentially conservative, approximation of the ModSel-CP set. However, in a regression setting where data points are continuously distributed, many canonical scores (e.g., the residual score, and other examples considered in Section 2.2) satisfy the additional conditions in the theorem, and consequently the upper bound in (11) differs from the ModSel-CP interval only by a set of measure zero. Consequently, in many common settings, the easy-to-calculate upper bound (11) is essentially an equivalent formulation for $\widehat{C}_{\textnormal{{ModSel-CP}}}(X_{n+1})$.
In other settings, the additional conditions of the theorem may not hold (e.g., due to ties among values). However, this is most likely to occur in discrete data settings, where $\mathcal{Y}$ is itself finite—i.e., a categorical response. In this setting, implementation is more straightforward since we can simply enumerate all possible values $y\in\mathcal{Y}$ and calculate the ModSel-CP prediction set directly—the upper and lower bounds are no longer necessary for computational efficiency.

### v3

# Conformal prediction after data-dependent model selection

@@ Abstract
Given a family of pretrained models and a hold-out set, how can we construct a valid conformal prediction set while selecting a model that minimizes the width of the set? If we use the same hold-out data set both to select a model (the model that yields the smallest conformal prediction sets) and then to construct a conformal prediction set based on that selected model, we suffer a loss of coverage due to selection bias. Alternatively, we could further splitting the data to perform selection and calibration separately, but this comes at a steep cost if the size of the dataset is limited. In this paper, we address the challenge of constructing a valid prediction set after data-dependent model selection—commonly, selecting the model that minimizes the width of the resulting prediction sets. Our novel methods can be implemented efficiently and admit finite-sample validity guarantees without invoking additional sample-splitting. We show that our methods yield prediction sets with asymptotically optimal width under certain notion of regularity for the model class. The improvement in the width of the prediction sets constructed by our methods are further demonstrated through applications to synthetic datasets in various settings and a real data example.
@@ 1 Introduction
However, despite the success of the conformal prediction framework in offering statistical validity, the optimality in the size of conformal prediction sets generally lacks a universal guarantee and often relies on a model-specific analysis with additional assumptions on the data distribution. For example, Sesia and Candès (2020) show that the conformalized quantile regression (Romano et al., 2019) achieves the optimal width asymptotically if the quantile regression function estimate is consistent and the data distribution has certain regularity properties. Similar results also hold for the split conformal method and the full conformal method when paired with the residual score (Lei et al., 2018). See also the work of Gyôrfi and Walk (2019); Burnaev and Vovk (2014); Sadinle et al. (2019); Izbicki et al. (2020), among others, for additional examples of settings where width optimality of conformal prediction has been established under specific conditions; in general, obtaining a small conformal prediction set is closely tied to choosing a conformity score that is well suited to the data distribution, as has been studied empirically in Aleksandrova and Chertov (2021); Johansson et al. (2017); Sesia and Candès (2020).
Given the necessity of choosing the right model to construct an effective prediction set, and the practical reality that the most suitable model cannot be predetermined without analyzing the data, an integrated framework that combines model selection with the construction of prediction sets would be highly beneficial. This serves as the motivation of this paper. In this work, we address the challenge of constructing a finite-sample valid conformal prediction set while selecting a model that minimizes the width of the prediction set, from a given family of pretrained models. Specifically, fix some target coverage level $1-\alpha\in(0,1)$, e.g., 90%. Given a collection of pretrained models, a set of independent and identically distributed hold-out data $\mathcal{D}_{n}$, and a new test point $X_{n+1}$ from the same distribution, we would like to construct a conformal prediction set $\widehat{C}_{n}$ using a selected model (i.e., selected as the model that offers the smallest prediction interval), such that
@@
  • the width of the prediction set $\widehat{C}_{n}$ is as small as possible (i.e., we have benefited from selecting the best model).
Several existing works in the literature have considered similar problems to the question described above. In Section 2.3 below, we will closely examine the work of Yang and Kuchibhotla (2024), which is most relevant to our work. In their work, Yang and Kuchibhotla (2024) propose either choosing the model with the shortest interval without correcting for selection to achieve small width (which suffers from selection bias unless the set of candidate models is extremely small), or using a sample splitting based approach (which removes the issue of selection bias, but suffers in statistical efficiency due to reduced sample size). In contrast, our work aims to correct for selection bias without any data splitting. We will compare our methods with theirs more closely below. For a more comprehensive discussion of related literature, we defer it to Section 6.1.
@@
We propose a novel method with two variants to construct finite-sample valid conformal prediction sets without additional sample-splitting after data-dependent model selection. Both versions are computationally efficient to implement, and offer validity with no assumptions aside from exchangeability of the data. In addition, we show that under certain regularity condition for the model class, our methods output prediction sets that are asymptotically optimal in width. We will see on simulated and real data that these methods yield tighter prediction sets in finite-sample relative to existing approaches, by removing selection bias without paying the price of data splitting.
@@
We first formulate our problem as well as surveying existing methods in Section 2. We then present our proposed methods as well as coverage guarantees in Section 3. The asymptotic optimality in width of our methods is investigated in Section 4. In Section 5, we conduct empirical comparisons across different methods to demonstrate the effectiveness of our methods in practice. Finally, we review the related literature and conclude with a brief discussion in Section 6.
@@ 2 Problem formulation
In this section, we first review some background on conformal prediction, which serves as the backbone of our prediction sets construction. We will then specify the class of model selection procedures considered in this paper and formulate the problem of interest. Finally, we will briefly review the existing methods related to our problem.
@@ 2.1 Background: conformal prediction
Consider the learning problem of predicting the response $y\in\mathcal{Y}$ based on an observed feature $x\in\mathcal{X}$. Conformal prediction offers a framework of constructing valid prediction sets without placing any assumption on the data distribution $P$ other than data exchangeability.
@@
If the amount of available data is limited, we may wish to avoid the cost of data splitting inherent in split conformal prediction (that is, in split conformal prediction, $n$ calibration data points have been removed from the total quantity of available data, and thus the score function $S$ is trained using a smaller sample size.) Full conformal prediction (Vovk et al., 2005) provides a solution that allows us to train a model and calibrate it to construct a prediction set without data splitting.
Given a dataset $\mathcal{D}_{n}\in(\mathcal{X}\times\mathcal{Y})^{n}$ and a symmetric algorithm $\mathcal{A}:\cup_{n\geq 0}(\mathcal{X}\times\mathcal{Y})^{n}\to\mathcal{F}$, where $\mathcal{F}$ is the set of measurable functions from $\mathcal{X}\times\mathcal{Y}$ to $\mathbb{R}$, the full conformal prediction set is defined as
$\widehat{C}_{\texttt{full-CP}}(X_{n+1})=\left\{y\in\mathcal{Y}:S^{y}(X_{n+1},y)\leq\mathrm{Quantile}_{1-\alpha}\left(\left\{S^{y}(X_{i},Y_{i})\right\}_{i\in[n]}\cup\{S^{y}(X_{n+1},y)\}\right)\right\},$ (2)
@@
2.2 The model selection criterion
@@
We next need to specify how we choose one of these pretrained models—what does it mean to choose the best, or most effective, model from the set? Formally, we will quantify model effectiveness through a loss function $\mathcal{L}$, as follows:
@@
A typical choice for this loss might be the average width of the prediction set,
@@
where ||$\cdot$|| ⋅ | denotes the size of the set, e.g., the Lebesgue measure if $\mathcal{Y}=\mathbb{R}$ (i.e., the width, if $C_{q}^{\lambda}$ returns intervals), or the cardinality if $\mathcal{Y}$ is a finite set of categorical labels. However, this is not the only choice we might use in practice—for instance, we might choose median, rather than mean, size. [^2] As an example of another kind of loss function that we might use in a regression setting, the work of Liang et al. (2023) consider a selection procedure that aims at minimizing the squared estimation error, $\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-f_{\lambda}(X_{i}))^{2}$, rather than the interval width. Like our work which will be described in details below, Liang et al. (2023)’s approach uses the full conformal framework to then incorporate a hypothesize test point $(X_{n+1},y)$ into this selection process. Note however that this requires a loss $\mathcal{L}$ that depends on $(X,Y)$ values at evaluation (rather than only $X$ values), and is therefore not directly comparable to our framework. See Section 6.1 for further discussion.
@@
    the width of the interval (which is shared across all values $x$, and therefore does not depend on the particular evaluation points $x_{1},\dots,x_{m}$).
@@
Main aim: Given a collection of pretrained models $\{S^{\lambda}:\lambda\in\Lambda\}$, and a calibration data set $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$, we aim to output a prediction set at a new test point $X_{n+1}$ that is both valid (under exchangeability of the data) and small (in the sense of aiming to minimize $\mathcal{L}$ over the collection of models).
@@ 2.3 Existing methods: the work of Yang and Kuchibhotla (2024)
YK-baseline.
In their first approach, split conformal prediction sets are constructed for each pretrained model $S^{\lambda}$, and the one with the smallest width is selected directly without any correction for selection bias. We refer this approach to as YK-baseline in this paper, since it provides a baseline against which other methods (that do account for selection bias) can be compared. We can define it precisely as follows. First, for each $\lambda\in\Lambda$, define
@@
In other words, this is the split conformal prediction set constructed with the selected model $\hat{\lambda}\in\Lambda$—but without any correction for selection bias. [^3] Yang and Kuchibhotla (2024)’s work does not formally specify a mechanism for defining a “best” prediction set, so here we choose a specific definition for concreteness, defining “best” to mean achieving the minimal loss (e.g., minimal width) over the $n+1$ calibration and test feature values.
@@
By construction, YK-baseline yields a prediction set with the smallest width among all the possible split conformal prediction sets $C_{\hat{q}(\lambda)}^{\lambda}$ (indexed over $\lambda\in\Lambda$). In terms of coverage, Yang and Kuchibhotla (2024) show the following concentration-based guarantee:
@@
Here their result assumes the data points $(X_{1},Y_{1}),\dots,(X_{n},Y_{n}),(X_{n+1},Y_{n+1})$ to be i.i.d., rather than only assuming exchangeability. We can interpret this lower bound as saying that, if $|\Lambda|$ is not too large (i.e., $\log(2|\Lambda|)\ll n$), then the selection bias is limited and will not substantially reduce coverage. On the other hand, for large $|\Lambda|$, this lower bound does not give a meaningful guarantee (i.e., coverage may be arbitrarily low due to selection bias), as will be confirmed in our experiments in Section 5.
YK-adjust.
@@
YK-split.
To avoid the problem of selection bias incurred by YK-baseline, Yang and Kuchibhotla (2024) propose an alternative method, based on data splitting, which we will refer to as YK-split. At a high level, the calibration set of $n$ data points is split into two parts, where $n_{1}$ data points are used to select a model, and the remaining $n_{2}=n-n_{1}$ data points are used to construct the split conformal prediction set with the selected model. Specifically, defining
@@ 3.1 The ModSel-CP method
As we will see in the proof, this marginal coverage result will follow by reformulating ModSel-CP as an instance of full conformal prediction. In particular, treating the model selection step (i.e., the process of selecting a model $\hat{\lambda}(y)$ as a function of the data) as a part of the training algorithm $\mathcal{A}$ in full conformal, the ModSel-CP method can be seen as conformalizing over a new meta-score output by this algorithm $\mathcal{A}$.
@@ 3.1.1 Implementation
For any given dataset $\mathcal{D}_{n}$ and test point $X_{n+1}$, the ModSel-CP prediction set defined in (9) satisfies
@@
Furthermore, if it holds almost surely that:
@@
Note that Theorem 3.2 does not provide an exact calculation for the prediction set $\widehat{C}_{\textnormal{{ModSel-CP}}}(X_{n+1})$, but rather provides upper and lower bounds on this set. Nevertheless, we argue that in most practical scenarios, the upper bound (11) can serve as a close approximation to the ModSel-CP set, as the additional conditions specified in the theorem are typically satisfied. To illustrate this, we consider the following example.
Theorem 3.2 in practice: the residual score.
In the case of the residual score $S^{\lambda}(x,y)=|y-f_{\lambda}(x)|$, where $f_{\lambda}$ is a pretrained regression function, the upper bound (11) is equal to a union of intervals,
$\widehat{C}_{\texttt{ModSel-CP}}(X_{n+1})\subseteq\left\{y\in\mathbb{R}:\min_{\lambda\in\mathcal{M}}|y-f_{\lambda}(X_{n+1})|\leq\hat{q}(\hat{\lambda})\right\}\\ =\bigcup\limits_{\lambda\in\left\{\lambda^{\prime}\in\Lambda:\hat{q}_{-}(\lambda^{\prime})\leq\hat{q}(\hat{\lambda})\right\}}\left[f_{\lambda}(X_{n+1})-\hat{q}(\hat{\lambda})\ ,\ f_{\lambda}(X_{n+1})+\hat{q}(\hat{\lambda})\right].$ (13)
In the meantime, the lower bound (12) is also equal to a union of intervals,
$\widehat{C}_{\texttt{ModSel-CP}}(X_{n+1})\supseteq\left\{y\in\mathbb{R}:\min_{\lambda\in\mathcal{M}_{-}}|y-f_{\lambda}(X_{n+1})|<\hat{q}(\hat{\lambda})\right\}\\ =\bigcup\limits_{\lambda\in\left\{\lambda^{\prime}\in\Lambda:\hat{q}_{-}(\lambda^{\prime})<\hat{q}(\hat{\lambda})\right\}}\left(f_{\lambda}(X_{n+1})-\hat{q}(\hat{\lambda})\ ,\ f_{\lambda}(X_{n+1})+\hat{q}(\hat{\lambda})\right).$ (14)
Since in most regression settings the data are continuously distributed, we have $\mathcal{M}=\mathcal{M}_{-}$ almost surely. Consequently, comparing (13) and (14), we see that the upper and lower bounds coincide up to a set of measure zero. Thus, the ModSel-CP set differs from the upper bound (13) only on a null set, effectively making the upper bound an alternative expression of the ModSel-CP set.
Returning to the general scores, we can equivalently rewrite the upper bound (11) as
@@
Hence, ModSel-CP — or more precisely, its upper bound used as a close approximation — can be efficiently implemented, provided that sets of the form $\left\{y:\mathcal{L}_{n+1}(\lambda,S^{\lambda}(X_{n+1},y))\leq q\right\}$ are easy to compute.
While the additional conditions in the theorem may occasionally fail to hold (e.g., due to ties among values), such cases mostly likely arise in discrete data settings, where $\mathcal{Y}$ is itself finite—i.e., a categorical response. However, this does not present a significant challenge to implementation in this case, since we can simply enumerate all possible values $y\in\mathcal{Y}$ and calculate the ModSel-CP prediction set directly—the upper and lower bounds are no longer necessary for computational efficiency.